1) Delia Landscaping is considering a new 4-year project. The necessary fixed assets will cost $187,000 and be depreciated on a 3-year MACRS and have no salvage value. The MACRS percentages each year are 33.33 percent, 44.45 percent, 14.81 percent, and 7.41 percent, respectively. The project will have annual sales of $124,000, variable costs of $33,100, and fixed costs of $12,650. The project will also require net working capital of $3,250 that will be returned at the end of the project. The company has a tax rate of 34 percent and the project's required return is 8 percent. What is the net present value of this project?
A)$35,840
B)$37,982
C)$38,845
D)$36,912
F)$32,853

2) Deep Mines has 43,800 shares of common stock outstanding with a beta of 1.54 and a market price of $51 a share. There are 10,000 shares of 7 percent preferred stock outstanding with a stated value of $100 per share and a market value of $83 a share. The 8 percent semiannual bonds have a face value of $1,000 and are selling at 96 percent of par. There are 5,000 bonds outstanding that mature in 13 years. The market risk premium is 7.5 percent, T-bills are yielding 3.6 percent, and the tax rate is 21 percent. What discount rate should the firm apply to a new project's cash flows if the project has the same risk as the company's typical project?
A)8.72%
B)9.30%
C)9.59%
D)9.17%
F)8.28%

Please show in details, not just answers thank you!
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

ANSWER : B : 37982

05:09 х 10 01 2020 Al166 ANSWER AB AC AF AL AA AD AE AG АН Al AJ AK 147 YEAR 3 COST OF FIXED ASSETS 148 -187000 149 WORKING C

ANSWER : B : 9.30%

05 19 A dx E ENG х 10 01 2020 fr 1200 A D Н 200 financial calculator strokes 201 FOR YTM first of all 202 BOND C/Y = P/Y = 20

05 19 A dx E ENG х 10 01 2020 fr 1249 D Н 224 225 226 CALCULATION OF WACC 227 NUMBER PRICE WEIGHTS COC(%) VEIGHTED 228 SOURCE